A 14-pound bowling ball feels a net force of -4.7 N as it slides down the lane. What is the acceleration of the bowling ball? (1 kg = 2.2 lbs)



1
Expert's answer
2021-03-02T18:04:50-0500

According the second Newton's law, the acceleration is:


"a = \dfrac{F}{m}"

where "F = -4.7N" is the net force, and "m = 14lbs = 14\/2.2kg" is the mass. Thus, obtain:


"a = \dfrac{-4.7\cdot 2.2}{14}\approx -0.74m/s^2"

Answer. "-0.74m\/s^2".